# Learning to Upsample and Upmix Audio in the Latent Domain

### v1

@@ 1 Introduction
Neural audio autoencoders have become central to modern audio technologies, adept at creating compact latent representations that preserve perceptually important information [1, 2, 3, 4]. These representations form the bedrock of many audio compression systems and enable powerful generative models, including text-to-audio synthesis via latent diffusion [5, 6, 7, 8, 9], next token prediction [10, 11, 12, 13, 14] and parallel decoding [15, 16, 17]. Despite the prevalence of autoencoders, many fundamental audio processing operations, such as bandwidth extension and mono-to-stereo upmixing, are still predominantly performed on raw waveform or spectral representations [18, 19, 20, 21].
@@
Figure 1: Top: Conventional post-generation/transmission workflow—decode to raw audio, then process it. Bottom: Our proposed method—perform all processing directly on the compact latent representation before the decoding step, avoiding costly operations on raw audio.
While prior work has explored audio processing within latent spaces, most methods employ computationally intensive large models. For example, bandwidth extension has primarily used diffusion-based approaches on continuous latents [23] or discrete tokens [24], and token-based transformers have been applied to source separation [25]. Although more computationally efficient source separation methods exist [26, 27, 28], they rely on high-dimensional intermediate representations that limit efficiency gains. Similarly, recent work repurposes neural audio codecs as mel-vocoders [29] or speech enhancers [30], but this requires fine-tuning the base autoencoder with expensive and complex time-domain objectives. In contrast, our approach tackles standard audio tasks using lightweight modules that operate directly on highly compressed latents produced by off-the-shelf, pre-trained audio autoencoders (achieving 16$\times$ or greater compression).
In this work, we demonstrate the effectiveness of our latent-domain processing framework through experiments on bandwidth extension and mono-to-stereo up-mixing. We demonstrate that operating entirely within the latent space allows for training models in less than two days on a single GPU. At inference, this approach achieves 100$\times$ efficiency gains while maintaining performance comparable to raw-audio post-processing techniques. This research establishes a more efficient and resource-conscious paradigm for audio processing, particularly beneficial for pipelines already incorporating neural audio autoencoders, thereby enabling faster, more integrated workflows across a variety of audio tasks. Crucially, the autoencoder-agnostic nature of our framework ensures that it is positioned to benefit from future advancements in audio autoencoder technology.
@@ 3.2 Training Details
Models are trained with purely latent space losses as described in Sec. 2 unless otherwise noted and in bfloat16 precision. For BWE where we use the discriminator, we set the following loss weights: $10.0$ for the $\mathcal{L}_{\mathrm{rec}}$ term, $0.5$ for $\mathcal{L}^{F}_{\mathrm{adv}}$, and $1.0$ for $\mathcal{L}^{F}_{\mathrm{fm}}$. While, when training the mono-to-stereo model we use: $10.0$ for $\mathcal{L}_{\mathrm{rec}}$ loss and $45e-$ 4 for the KL loss term. Models are trained with a batch size of $256$ on audio chunks with a duration of 1.4 s for BWE, and 4 s for mono-to-stereo. We perform 250K training steps on a single H100-80GB GPU. We use AdamW with a learning rate of $45e-$ 4 for the main model and $41e-$ 4 for the latent discriminator, we also use linear warm-up totaling 1K steps for the main model and 20K steps for the latent discriminator. Weight decay is set at $41e-$ 4.
@@ 3.3.1 Baselines
For each task, we select specialized state-of-the-art (SOTA) baselines for comparison: MusicHiFi-BWE [19] and Aero [18] for bandwidth extension, and MusicHiFi-M2S [19] for mono-to-stereo comparisons. In every task, we evaluate two scenarios: one where the baseline model has access to the undistorted input, and another involving autoencoder transmission where we first pass the model input through the autoencoder (denoted as $\mathrm{VAE+Baseline}$).
@@ 4.2 Mono-to-Stereo
Table 2 row 2 shows the raw audio baseline’s performance (0.85 STFT-D) without a transmission constraint. Row 3 presents the baseline performance with a transmission constraint (1.34 STFT-D left/right), operating on the decoded mono VAE output. MusicHiFi copies the mono input to the center and estimates a side channel. In contrast, our method estimates the left/right VAE latents directly, since side can be slightly out-of-domain for the VAE. We first evaluate our module by sampling the control vector from the prior, forgoing control. As shown in row 4, this latent module (1.31 STFT-D left/right) surpasses the transmitted raw audio baseline (1.34 STFT-D left/right) in left/right error. To demonstrate the potential of control, we provide the latent module with the oracle control vector, derived from the ground-truth stereo. In row 5 (1.22 STFT-D left/right), the controlled latent module improves performance for all metrics and formats with the exception of mel-D on side. Comparing computational costs, our latent module requires 138×138$\times$138 × fewer FLOPS than MusicHiFi.
@@ 5 Conclusion
In this work, we introduced the Re-Encoder framework, an approach for performing audio processing operations entirely within the latent space of pre-trained neural audio autoencoders. This directly addresses the inherent inefficiency of operating on high-dimensional raw audio representations for audio-to-audio tasks. By leveraging these compact latents, our method dramatically simplifies the training process. It requires only a straightforward latent $L_{1}$ reconstruction loss, optionally augmented by a single latent adversarial discriminator or latent condition encoder. This stands in stark contrast to the complex multi-scale loss and multiple discriminator setups typically needed for conventional raw-audio methods. Through experiments in bandwidth extension and mono-to-stereo up-mixing, we demonstrated significant efficiency gains, achieving up to a 200×200$\times$200 × speedup while maintaining output quality. By keeping operations entirely within the efficient latent domain, the Re-Encoder establishes a more efficient paradigm for audio processing pipelines which use autoencoders during both training and inference. This paves the way for faster, more resource-efficient workflows across a variety of audio tasks.

### v2

@@ 1 Introduction
Neural audio autoencoders have become central to modern audio technologies, adept at creating compact latent representations that preserve perceptually important information [1, 2, 3, 4]. These representations form the bedrock of many audio compression systems and enable powerful generative models, including text-to-audio synthesis via latent diffusion [5, 6, 7, 8, 9, 10], next token prediction [11, 12, 13, 14, 15] and parallel decoding [16, 17, 18]. Despite the prevalence of autoencoders, many fundamental audio processing operations, such as bandwidth extension and mono-to-stereo upmixing, are still predominantly performed on raw waveform or spectral representations [19, 20, 21, 22, 23].
@@
Figure 1: Bottom: Conventional post-generation/transmission workflow—decode to raw audio, then process it. Training involves multiple reconstruction and adversarial objectives. Top: Our proposed method—perform all processing directly on the compact latent representation before the decoding step, avoiding costly operations on raw audio. Training performed solely in the latent domain.
[…]
@@ 3.2 Training Details
Models are trained with purely latent space losses as described in Sec. 2 unless otherwise noted and in bfloat16 precision. For BWE where we use the discriminator, we set the following loss weights: $10.0$ for the $\mathcal{L}_{\mathrm{rec}}$ term, $0.5$ for $\mathcal{L}^{F}_{\mathrm{adv}}$, and $1.0$ for $\mathcal{L}^{F}_{\mathrm{fm}}$. While, when training the mono-to-stereo model we use: $10.0$ for $\mathcal{L}_{\mathrm{rec}}$ loss and 5e45e-4 for the KL loss term. Models are trained with a batch size of $256$ on audio chunks with a duration of 1.4 s for BWE, and 4 s for mono-to-stereo. We perform 250K training steps on a single H100-80GB GPU. We use AdamW with a learning rate of 5e45e-4 for the main model and 1e41e-4 for the latent discriminator, we also use linear warm-up totaling 1K steps for the main model and 20K steps for the latent discriminator. Weight decay is set at 1e41e-4.
@@ 3.3.1 Baselines
For each task, we select specialized state-of-the-art (SOTA) baselines for comparison: MusicHiFi-BWE [20] and Aero [19] which outperform models like AudioSR [25, 20] for bandwidth extension, and MusicHiFi-M2S [20] for mono-to-stereo comparisons. In every task, we evaluate two scenarios: one where the baseline model has access to the undistorted input, and another involving autoencoder transmission where we first pass the model input through the autoencoder (denoted as $\mathrm{VAE+Baseline}$).
@@ 4.2 Mono-to-Stereo
Table 2 row 2 shows the raw audio baseline’s performance (0.85 STFT-D) without a transmission constraint. Row 3 presents the baseline performance with a transmission constraint (1.34 STFT-D left/right), operating on the decoded mono VAE output. MusicHiFi copies the mono input to the center and estimates a side channel. In contrast, our method estimates the left/right VAE latents directly, since side can be slightly out-of-domain for the VAE. We first evaluate our module by sampling the control vector from the prior, forgoing control. As shown in row 4, this latent module (1.31 STFT-D left/right) surpasses the transmitted raw audio baseline (1.34 STFT-D left/right) in left/right error. To demonstrate the potential of control, we provide the latent module with the oracle control vector, derived from the ground-truth stereo. In row 5 (1.22 STFT-D left/right), the controlled latent module improves performance for all metrics and formats with the exception of mel-D on side. Comparing computational costs, our latent module requires $138\times$ fewer FLOPS than MusicHiFi.
@@ 5 Conclusion
In this work, we introduced the Re-Encoder framework, an approach for performing audio processing operations entirely within the latent space of pre-trained neural audio autoencoders. This directly addresses the inherent inefficiency of operating on high-dimensional raw audio representations for audio-to-audio tasks. By leveraging these compact latents, our method dramatically simplifies the training process. It requires only a straightforward latent $L_{1}$ reconstruction loss, optionally augmented by a single latent adversarial discriminator or latent condition encoder. This stands in stark contrast to the complex multi-scale loss and multiple discriminator setups typically needed for conventional raw-audio methods. Through experiments in bandwidth extension and mono-to-stereo up-mixing, we demonstrated significant efficiency gains, achieving up to a $200\times$ speedup while maintaining output quality. By keeping operations entirely within the efficient latent domain, the Re-Encoder establishes a more efficient paradigm for audio processing pipelines which use autoencoders during both training and inference. This paves the way for faster, more resource-efficient workflows across a variety of audio tasks.